4x^2+3y^2+z^2-8xy-6yz+4zx


1
Expert's answer
2021-08-19T07:04:22-0400

Solution.

Write the given form as "4x^2-4xy+2xz-4xy+3y^2-3yz+2zx-3yz+z^2"

The corresponding matrix is "A=\begin{pmatrix}
 4 & -4&2 \\
 -4 & 3&-3\\
2&-3&1
\end{pmatrix}"

We write "A=IAI"

We apply elementary operations on A and we apply the same row operations on the prefactor (R) and column operations on the post factor (C).

"\begin{pmatrix}
 4 & -4&2 \\
 -4 & 3&-3\\
2&-3&1
\end{pmatrix}=\begin{pmatrix}
 1 & 0&0 \\
 0&1&1\\
0&0&1
\end{pmatrix}A\begin{pmatrix}
 1 & 0&0 \\
 0&1&1\\
0&0&1
\end{pmatrix}"

Applying "R_2+R_1, 2R_3-R_1, 2R_2-R_3,R_1-2R_3,R_1+R_2,R_1/2,
R_2/\sqrt{2}, R_3/\sqrt{2},R_2\leftrightarrow R_3"

and "C_2+C_1, 2C_3-C_1, 2C_2-C_3, C_1-2C_3, C_1+C_2,C_1/2,
C_2/\sqrt2,C_3/\sqrt2,C_2\leftrightarrow C_3"

we get

"\begin{pmatrix}
 1 & 0&1 \\
 0 & -1&0\\
0&0&-1
\end{pmatrix}=\begin{pmatrix}
 3& 1&-3 \\
 -1/\sqrt{2}&0&2/\sqrt2\\
3/\sqrt2&2/\sqrt2&-2/\sqrt2
\end{pmatrix}A\begin{pmatrix}
 3& -1/\sqrt2&3/\sqrt2 \\
 1&0&2/\sqrt2\\
-3&2/\sqrt2&-2/\sqrt2
\end{pmatrix}"

We have "D=P^TAP," where "D=\begin{pmatrix}
 1 & 0&1 \\
 0 & -1&0\\
0&0&-1
\end{pmatrix}."

"P=\begin{pmatrix}
 3& -1/\sqrt2&3/\sqrt2 \\
 1&0&2/\sqrt2\\
-3&2/\sqrt2&-2/\sqrt2
\end{pmatrix}."

The linear transformation is X=PY, or

"\begin{pmatrix}
 x \\
 y\\z
\end{pmatrix}=\begin{pmatrix}
 3& -1/\sqrt2&3/\sqrt2 \\
 1&0&2/\sqrt2\\
-3&2/\sqrt2&-2/\sqrt2
\end{pmatrix}\begin{pmatrix}
 y_1\\
 y_2\\y_3
\end{pmatrix}"

From here


"x=3y_1-1/\sqrt2 y_2+3/\sqrt2y_3,\newline
y=y_1+\sqrt2 y_3,\newline
z=-3y_1+\sqrt2 y_2-\sqrt2 y_3."


The canonical form is reduced to "X^TAX=Y^TDY" .

We will have

"\begin{pmatrix}
 y_1&y_2&y_3 \\
\end{pmatrix}\begin{pmatrix}
 1 & 0&1 \\
 0 & -1&0\\
0&0&-1
\end{pmatrix}\begin{pmatrix}
 y_1 \\
 y_2\\y_3
\end{pmatrix}=y_1^2-y_2^2-y_3^2."

Answer. The given quadratic form is reduced to normal form "y_1^2-y_2^2-y_3^2."